OpenStudy (anonymous):

Which shows x^2+9 factored over the set of complex numbers? a. (x+3i)^2 b. (x+3i)(x-3i) c. (x-1)(x+9i) d. (-x+3i)(x-3i)

OpenStudy (inkyvoyd):

HINT: (a+b)(a-b)=a^2-b^2

OpenStudy (anonymous):

Okay! Thanks, I'm going to try that

OpenStudy (anonymous):

I think the answer is A. Is that correct?

OpenStudy (inkyvoyd):

No.

OpenStudy (anonymous):

Okay, then I think its b

OpenStudy (inkyvoyd):

Why?

OpenStudy (anonymous):

because (x+3i)(x-3i)= x^2+9

OpenStudy (inkyvoyd):

Did you expand it out?

OpenStudy (anonymous):

No I didn't expand it.

OpenStudy (inkyvoyd):

You'll never know the answer.

OpenStudy (anonymous):

Dude, are you gonna help me or not?

OpenStudy (anonymous):

do you know how to expand it out?

OpenStudy (inkyvoyd):

I'm not going to tell you the answer, but I will help you if you are willing to put forth effort.

OpenStudy (anonymous):

If you don't know how to expand it out here is an example \[(x+3i)^2\] (x+3i)(x+3i) x^2 +3i +3i +9i^2 (i^2 is 1 so its just 9*1) \[x^2 +6i +9 \]

OpenStudy (anonymous):

my mistake i^2 is actually -1 so the answer would be \[x^2 +6i -9\]

OpenStudy (anonymous):

All in all your answer is B if you never figured it out